User Tools

Site Tools


lbaops:lbafeb2008:todo

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
lbaops:lbafeb2008:todo [2008/02/03 23:03]
chris.phillips
lbaops:lbafeb2008:todo [2008/02/04 09:44]
chris.phillips
Line 1: Line 1:
-**Parkes**\\\ +**Parkes**\
-\\\ +
-  *  //Test Apple Xserve RAIDS// **DONE**\\\ +
-  *  Connect DAS to pkvsi1 (not v1pks)\\\ +
-  *  Check VSIC connections\\\ +
-  *  Test DAS \\\ +
-\\\ +
-**Mopra**\\\ +
-\\\ +
-  *  //Copy DAS profiles from v1mop// **DONE**\\\ +
-  *  Reconfig network setup\\ +
-  *  Get turret rotation/​translation working again\\\ +
-  *  Verify 12mm Rx setup/​operation\\\ +
-  *  Verify all updated software working.\\\ +
-  *  Setup hybrid for circular polarisation at 12mm. May have to be done on Wedn.\\\ +
-\\\ +
-**ATCA**'​\\\ +
-\\\ +
-  *  //Reconnect VSIC and test// **DONE**\\\ +
-  *  //Test dodgey channel on VSIC/​cables//​ **DONE**\\\ +
-\\\+
 \ \
-**General**\\\ +  ​ //​Test Apple Xserve RAIDS// ​**DONE**\ 
-\\\ +  *  Connect DAS to pkvsi1 (not v1pks)\ 
-  *  Create VC fringes test experiments\\+  *  Check VSIC connections
-  *  Check CVS versio upto date\\+  ​* ​ Test DAS \ 
-  *  Merge UDP version?\\+\ 
-  *  Check cdisko working OK\\+**Mopra**\ 
-  *  Investigate cdisko scheduling\\+
-  *  Check RtFC works with green\\+  *  //Copy DAS profiles from v1mop// **DONE**\ 
-  *  Copy schedules to all observatories and run atcasched at Mopra/ATCA\\+  *  //Reconfig network setup// **DONE**\ 
-  *  Label all disks\\\+  *  Get turret rotation/​translation working again\ 
 +  *  Verify 12mm Rx setup/​operation\ 
 +  *  Verify all updated software working.\ 
 +  *  Setup hybrid for circular polarisation at 12mm. May have to be done on Wedn.\ 
 +
 +**ATCA**'​\ 
 +
 +  *  //Reconnect VSIC and test// **DONE**\ 
 +  *  //Test dodgey channel on VSIC/​cables//​ **DONE**\ 
 +
 +**Hobart**\ 
 +
 +  *  Check CVS version\ 
 +
 +**Ceduna**\ 
 +
 +  *  Check CVS version\ 
 +  *  Check can connect to green\ 
 +
 +
 +**General**\ 
 +
 +  *  Create VC fringes test experiments\ 
 +  *  Check CVS versio upto date\ 
 +  *  Merge UDP version?\ 
 +  *  Check cdisko working OK\ 
 +  *  Investigate cdisko scheduling\ 
 +  *  Check RtFC works with green\ 
 +  *  Copy schedules to all observatories and run atcasched at Mopra/​ATCA\ 
 +  *  Label all disks\
   *  Update the disk status page on the wiki with disk IDs   *  Update the disk status page on the wiki with disk IDs
  
lbaops/lbafeb2008/todo.txt · Last modified: 2015/12/18 16:38 (external edit)